STAR Protected
December 22
- RAW to MuDst production
- pp500_production_2017 (physics)
389582 files have been converted out of 654286
Around 60% Completion
- pp500_production_2017 (physics)
FST QC meeting
virtual , at 14:00 (GMT), duration : 00:00
Internal review of QC process and current status
Recommendation from review
An internal review of the procedures and implementation of the QC plans and documentation should be performed after the first 10% of the production is complete before further expensive and sparse parts (sensors and asics!) are committed.
The brief review should present
1) Demonstrate implementation of e-log
Take 2 cases
a) One of the prototype that was populated and tested (03,/04) Try to grade it with the current requirements as we know they were successfully populated and running
b) Take of the currently produced modules that were send to UIC (9-12) and show the grading
2) Present a summary of grading for the 09- now modules for each of the sub-grading quantities (outer, inner, different fastnesses, DR). Explain how the weighted grading is done.
3) Present what went wrong with the rejected ones (overflow glue,..) and what that implies for further production
Charge to review:
Is the QC documentation sufficient?
Does the QC documentation in the assembly document (V3) reflect what is done now?
Should the summary contain some of the sub grading?
Is there a need for any change in procedures, or QC measures?
.
Journal Review
Here is the material form the PLB review
1. Nov. 7 replies and draft changes
Paper Draft
- In V1 Fig. 6 was added
- In V2 Fig. 6 and 7 reduced in size
Analysis Note
Added section on z-vertex efficiency in V1
In V3 updated tables in the Appendix, also Fig. 21 updated.
tpc weekly
virtual , at 20:00 (GMT), duration : 00:00
-- hardware
-- software
Minutes
TPC weekly meeting
Hardware; Alexei excused. Is working on laser and gas system
The GG driver modules have been installed (Christian, as Tim is out) One module does not provide HV. There are two spare modules.
David is communicating with Tonko how best to provide the SC for the Gating Grid.
Software
Irakli had at LBNL meeting; provided an update to me after meeting
“I talked to Yuri last week about progress and the calculation of the distortions.
I also pointed out to him a little problem in the reading of the inner sector position parameters from the database. We reached out to Dimitry and he changed the readout for that table. It hopefully was not affecting anything, but now I have a good way to compare apples to apples as far as the inner sector position parameters are concerned.
While doing so I noticed that that is where a huge shift in z direction originates, so now I want to see if that is the right result or an artifact from what I've been saying – one can end up in a local minima for the alignment procedure. I will keep you guys updated I the days to come. I am running these jobs as fast as possible (Dmitry's fix has been applied only Monday).”
Yuri presented some intermedial results on his investigations of showing effect of his modified equations as discussed in the previous meeting 12/9/2020 that also have the documents
https://www.star.bnl.gov/~fisyak/star/Tpc/Distortions/
Unless there are updates we need to discuss in person this was the last TPC meeting for 2020. Please send updates to the mailing list until then 1/5/2021
zoom info
December 15
- RAW to MuDst production
- pp500_production_2017 (physics)
388055 files have been converted out of 654286
Around 60% Completion
- pp500_production_2017 (physics)
BTof issue in Run12 Cu+Au production
TPC weekly
zoom meeting, at 20:00 (GMT), duration : 00:00
Agenda:
-hardware status
- software
--- alignment progress
--- https://drupal.star.bnl.gov/STAR/system/files/he34embedding4th.pdf
--- AOB
TPC meeting minutes December 9, 2020
Hardware:
Alexei is done with laser maintenance on West side; waiting for etof re-install to be complete on East (done)).
Bob and Tonko finishing the work on Westside, which will close first. One more slicing to be done Wed. Still 5 splicing to be done on East side
Irakli have reported to Yuri he is revisiting the alignment (inner-outer) from the cosmics.
Yue-Hang gave update on 3He eff . Guannan had looked at effect on RMS ped in Sl10k in embedding and see no effect. In never libraries the RMSX cut look ok excepts for eta~0 .
At eta ~0 there may be other issues- as seen for other particles. Yuri points out there may be influence of prompt there |z| > 135, though issues are seen at up to ~20cm (??) from readout chambers.
Yuri reported on work on space charge correction and reports there could be a difference between East and West corrections. It is work in progress. Initial report is posted on meeting page, as well as some note by Jim Thomas from ALICE work on this subject.
Time | Talk | Presenter |
---|---|---|
15:20 | TPC Distortions (progress report) ( 00:20 ) 2 files | Yuri Fisyak |
December 8
- RAW to MuDst production
- pp500_production_2017 (physics)
386585 files have been converted out of 654286
Around 59% Completion
- pp500_production_2017 (physics)
December 1
- RAW to MuDst production
- pp500_production_2017 (physics)
384172 files have been converted out of 654286
Around 59% Completion
- pp500_production_2017 (physics)
Cold QCD/ Spin PWG Meeting
https://drupal.star.bnl.gov/STAR/pwg/spin/weekly-pwg-meetings, at 00:00 (GMT), duration : 00:00
Cold QCD/ Spin PWG Meeting
https://drupal.star.bnl.gov/STAR/pwg/spin/weekly-pwg-meetings, at 00:00 (GMT), duration : 00:00
Cold QCD/ Spin PWG Meeting
https://drupal.star.bnl.gov/STAR/pwg/spin/weekly-pwg-meetings, at 00:00 (GMT), duration : 00:00
Cold QCD/ Spin PWG Meeting
https://drupal.star.bnl.gov/STAR/pwg/spin/weekly-pwg-meetings, at 00:00 (GMT), duration : 00:00
TPC weekly
virtual , at 20:00 (GMT), duration : 00:00
Minutes:
Tim reported on Gating Grid status
Crate has all modified modules. It was tested locally.
The web based script by Tonko does not seem to work after the upgrades. He is aware of it.
The polarity is swtiched at the connector for all.
The sector 8 is special and should be checked.
FV will contact David on the slow control of the gating grid; how should it be implemented?
Alexei has been working on laser system
- optics replaced; signal looks good
- started with alignment of akll rafters
- propbably will need another week
Software
- no discussion as some key people could not join the meeting today. There was a longer discussion on the run-19 alignment at the monthly S&C meeting
- It is of high priority to get this resolved; will have as agenda point next week